OpenStudy (anonymous):

Triple integral of (4*cos(x^2))/(2*sqrt(z)), given that 2*y<=x<=2, 0<=y<=1, 1<=z<=4. I know that you have to use an infinite series for cos(x^2), and according to Maple the answer is sin(4). I can't seem to get that answer!

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\int\limits_{1}^{4}\int\limits_{2y}^{2}(2\cos (x^2))/(\sqrt{z})dx dz dy\]

OpenStudy (inkyvoyd):

I"m going to have to give you a medal simply because that math looks so esoteric from where I'm at.